OpenStudy (anonymous):

Prove sqrt(3) is irrational.

OpenStudy (saifoo.khan):

it doesnt have real roots.

OpenStudy (anonymous):

how will u prove it, without using calculator

OpenStudy (anonymous):

Also, there is an actual proof...

OpenStudy (anonymous):

ye, and thats what

OpenStudy (anonymous):

Okay, but I demand a fanship and a medal, because it's tedious to write.

OpenStudy (anonymous):

lol nevermind :P

OpenStudy (anonymous):

WELL FINE THEN. :P

OpenStudy (anonymous):

you want an actual proof right?

OpenStudy (anonymous):

as usual with such proofs start with assuming the contrary, that \[\sqrt{3}=\frac{a}{b}\] with a, b integers. the square and get \[3=\frac{a^2}{b^2}\iff 3b^2=a^2\]

OpenStudy (anonymous):

I was going to give him Euclid's Proof too, but he didn't want it....

OpenStudy (anonymous):

now appeal to the fundamental theorem of arithmetic, and factor a and b as the product of primes.

OpenStudy (anonymous):

oh then i will be quiet